if the base of a rectangle is increased by 30% and the altitude is decreased by 20% the area is increased by
Answer by DrBeeee(684)

(1) A = b*a
The base is increased by 30% or the new base is 1.30*b, and the altitude is decreased by 20% or the new altitude is 0.80*a
New area, An is given by
(2) An = (1.3*b)*(0.8*a) or
(3) An = 1.04*b*a or
(4) An = (1 + .04)*b*h or
(5) An = b*a + 0.04*b*a or
(6) An = A + 4% of A
Answer: The area is increased by 4%